A firm is considering to undertake one of these four projects. Recruit new staff. This would generate GH¢ 1,310 profit to the firm. Buy raw materials; which would generate GH¢ 1,420 profit. Pay workers’ bonuses. This would earn the firm GH¢ 1,500. Open another branch; which would make the firm record GH¢ 1,480 as profit. Each of the projects will cost the firm GH¢ 1,000.

(a) Rank the projects in order of importance, from the first to the last

(b) Find the opportunity cost of each project of the first three items in (a). (c) Find the net benefit of each of the first three items in (b) using

(i) direct cost.

(ii) opportunity cost.

(d) From your results in (cii), which project will you recommend to the firm? Explain your answer.


1
Expert's answer
2022-02-16T08:28:46-0500

(a) Rank the projects in order of importance, from the first to the last

  1. Pay workers’ bonuses. This would earn the firm GH¢ 1,500.
  2. Open another branch; which would make the firm record GH¢ 1,480 as profit
  3. Buy raw materials; which would generate GH¢ 1,420 profit.

(b) Find the opportunity cost of each project of the first three items in (a).

  • The opportunity cost of paying workers to opening another branch is:

Opportunity cost = "\frac{cost\, of\, paying\, workers}{cost\, of\, opening\,branch}" = "\frac{1500}{1480}" = GH¢. 1.014

  • The opportunity cost of opening another branch to paying workers is:

Opportunity cost = "\frac{cost\, of\, opening\,branch}{cost\, of\, paying\, workers}" = "\frac{1480}{1500}" = GH¢. 0.987

  • The opportunity cost of opening another branch to buying raw materials is:

Opportunity cost = "\frac{cost\, of\, opening\,branch}{cost\, of\, raw\,materials}" = "\frac{1480}{1420}" = GH¢. 1.042

  • The opportunity cost of buying raw materials to opening another branch is:

Opportunity cost = "\frac{cost\, of\, raw\,materials}{cost\, of\, opening\,branch}" = "\frac{1420}{1480}" = GH¢. 0.959

  • The opportunity cost of buying raw materials to paying workers is:

Opportunity cost = "\frac{cost\, of\, raw\,materials}{cost\, of\, paying\, workers}" = "\frac{1420}{1500}" = GH¢. 0.947

  • The opportunity cost of buying raw materials to paying workers is:

Opportunity cost = "\frac{cost\, of\, paying\, workers}{cost\, of\, raw\,materials}" = "\frac{1500}{1420}" = GH¢. 1.056

 (c) Find the net benefit of each of the first three items in (b) using

(i) direct cost.

Net benefit = total benefit - total cost

= (1500 + 1480 + 1420) - (1000 + 1000 + 1000)

Thus Net benefit = GH¢ 1400


(ii) opportunity cost.


Opportunity Cost of paying workers

Opportunity Cost of paying workers = sum of Explicit Cost (paying workers) + Implicit Cost (forgone profits from opening branch) to Pursue.

Opportunity Cost = 1000 + 480 = GH¢ 1480

Opportunity Cost of paying workers = sum of Explicit Cost (paying workers) + Implicit Cost (forgone profits from raw materials) to Pursue.

Opportunity Cost = 1000 + 420 = GH¢ 1420


Opportunity Cost of opening a new branch

Opportunity Cost of opening new branch = sum of Explicit Cost (opening new branch) + Implicit Cost (forgone profits from paying workers) to Pursue.

Opportunity Cost = 1000 + 500 = GH¢ 1500

Opportunity Cost of opening new branch = sum of Explicit Cost (opening new branch) + Implicit Cost (forgone profits from raw materials) to Pursue.

Opportunity Cost = 1000 + 420 = GH¢ 1420


Opportunity Cost of raw materials

Opportunity Cost of raw materials = sum of Explicit Cost (raw materials) + Implicit Cost (forgone profits from paying workers) to Pursue.

Opportunity Cost = 1000 + 500 = GH¢ 1500

Opportunity Cost of raw materials = sum of Explicit Cost (raw materials) + Implicit Cost (forgone profits from opening a new branch) to Pursue.

Opportunity Cost = 1000 + 480 = GH¢ 1480


(d) From your results in (c. ii), which project will you recommend to the firm? Explain your answer.

If they are interested in profit, they should invest in paying workers.

Because the economic profit benefit of paying workers is 1500 - 1420 = GH¢. 80

Which is the highest economic profit among the three.